OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. B(6, 2), m =

OpenStudy (anonymous):

y = 2 , x = 6 m = y2-y1/ x2-x1 y = m (x1-x2) 2 = m (x - 6 ) 2 = mx - 6m am stuck ><" !

OpenStudy (anonymous):

how do you write it in standard form

OpenStudy (anonymous):

He didn't give the slope, so there's actually nothing we can do.

OpenStudy (anonymous):

the standard form is y = mx + b

OpenStudy (anonymous):

tropical33 , i see thats y am stuck there xD !

OpenStudy (anonymous):

B(6, 2), m = -1/2

OpenStudy (anonymous):

2=6*-0.5 + b 2=-3+b b=5 so: y=-0.5x+5

OpenStudy (anonymous):

in standard form